# From CVE Entries to Verifiable Exploits: An Automated Multi-Agent Framework for Reproducing CVEs

### v1

@@ 4.1.2 Exploiter
Candidate Selection. Generating a working exploit from a given CVE in a pre-configured vulnerable environment is a crucial task for reproducing CVEs. However, the Exploiter requires less code base exploration than the Builder, as advisories often provide PoCs, exploitation steps, or vulnerability details. Thus we include expensive reasoning models like o3, claude-3.7-sonnet, and gemini-2.5-pro to the candidates list of developers. To mitigate instability seen in open-source models (Section 4.1.1), we employ only closed-source LLMs (Table 3) as developers, while retaining the same critic LLMs as in the Builder.
Developer Evaluation. For the 32 successful setups from Section 4.1.1, we ran CVE-Genie up to the Exploit Developer stage 6 and manually evaluated each developer LLM as follows: (1) we checked whether the LLM executed a PoC or command sequence that demonstrated the exploit, (2) we verified that the vulnerability was clearly triggered (e.g., uncontrolled recursion in CWE-674 resulting in a Maximum Recursion Error), (3) we assessed fidelity to the CVE description, (4) we ensured no mock-up environments or fake exploits were used, and (5) we confirmed the exploit script followed the expected format (see Section 3.3.1).
Among all models, o3 consistently performed best, successfully reproducing the highest number of exploits (13 out of 27) and demonstrating them end-to-end within the actual vulnerable environment. In contrast, o4-mini often produced incorrect or mock-up exploits, and struggled with runtime errors. While claude and gemini models showed strong capabilities in demonstrating exploits, they lacked versatility across different vulnerability types. Hence, we select o3 as the most effective Exploit Developer.
Critic Prompt Design. We examined the failed exploit attempts and categorized the recurring error patterns. These error types directly mirrored our manual verification criteria (e.g., failure to demonstrate the exploit, unclear triggering of the vulnerability, deviation from the CVE description, reliance on dummy/fake exploits, or incorrect script formatting). We then incorporated these categories into the critic agent’s prompt, allowing it to automatically detect the same issues in the Exploit Developer logs.
@@
Feedback Assessment. We evaluated feedback effectiveness by running CVE-Genie on 32 correct setups with the top developer (o3) and critic (o4-mini). This yielded valid exploits for 16 CVEs, of which 3 were fixed based on the critic’s feedback. As in Section 4.1.1, the developer mainly resolved minor issues (e.g., incomplete verification or missing log triggers), while more complex flaws (e.g., incorrect exploit verification) persisted. To balance usefulness and cost, we therefore restrict feedback to a single iteration, focusing on quick corrections of minor errors.
@@ 4.1.3 Verifier
Critic Evaluation. As shown in Figure 7(b), only the o3 model successfully identified critical but subtle issues in the verifier scripts, such as weak or flawed verification logic (Figure 4). Other models provided weak or no critique, making o3 the most effective critic for the CTF Verifier.
Feedback Assessment. We evaluated CVE-Genie on 16 correct exploits from the Exploiter, successfully generating accurate verifiers for 15 of them, of which 4 verifiers were improved through critic feedback. In our experiment, the maximum number of feedback iterations required to resolve either functional or critique-based issues was 4; therefore, we allow up to 5 retries for both the Flag Checker and the Verifier Critic.
@@ 4.1.4 Time and Cost Constraints
4.2 Diversity-Aware Baseline Assessment
@@
Dataset. For this study we use the same dataset $D_{S}$ of 60 diverse CVEs (see Section 4.1).
Methodology. Due to the non-deterministic nature of LLMs, a single run may fail due to incorrect reasoning, while later attempts might succeed [33]. To account for this, we run CVE-Genie multiple times on the CVEs in $D_{S}$, using its most optimal configuration (Section 4.1). After each iteration, we store successfully reproduced CVEs and rerun on the remaining ones. This iterative approach helps maximize CVE reproduction and reveals CVE-Genie’s convergence over seven runs.
Results. We use statistical methods as well as manual analysis of randomly selected 25 CVE reproduction runs, and we categories our results and analysis in the following categories:
Variance Over Multiple Runs. The variability in CVE-Genie’s performance primarily stems from the Builder phase, which is more open-ended than the targeted exploit generation guided by CVE context. Project setup involves repository exploration (Pre-Requisite Developer) and command execution (Setup Developer), where environment-specific behaviors often cause failures. For example, one run failed because the Setup Developer modified the PATH variable, while a clean retry succeeded, highlighting the value of reattempting in a fresh environment. Since the CTF Verifier depends on the Exploiter, which in turn relies on the Builder, this initial project setup is a critical bottleneck. As a previously failed build, if later successful, can enable downstream modules to reproduce a new CVE (as illustrated in Figure 8). Over multiple runs, the number of successful project, exploit, and verifier builds showed consistent gains with convergance at run seven, indicating that most recoverable failures can be resolved within a few retries.
Vulnerability and Project Context. Across 60 CVE reproduction attempts, we observed a 63.3% (38/60) overall success rate, with PoC availability in the advisory significantly helping the reproducibility (60% with PoC). Web-centric ecosystems (TypeScript/JavaScript/PHP) showed markedly higher success than compiled/system-heavy stacks (C/C++/Go/C#/Kotlin/Java), where build and environment issues dominated failures. Classic input-validation and injection classes (e.g., SQLi, XSS, path traversal, prototype pollution) reproduced reliably when endpoints were reachable, whereas access-control vulnerabilities underperformed, likely due to the need for realistic authentication/authorization context. The principal blockers were build failures (36%) and timeouts (40%), indicating that improvements in environment provisioning and adaptive budgeting are as critical as exploit logic for scaling automated CVE reproduction.
RQ1: CVE-Genie’s success first depends on whether it can build the vulnerable project, after which performance varies with vulnerability type and language, e.g., web-based and input-validation bugs are easier to reproduce, while compiled/system-heavy stacks and access-control related CVEs are harder.
@@ 4.3 Robustness to Loss of CVE Context
Dataset ($D_{R}$). To evaluate CVE-Genie’s robustness to incomplete CVE data, we perform an ablation study using 15 CVEs from $D_{S}$, each with complete information (i.e., CVE description, PoC, patch commit, and security advisory) and reproduced in the first three baseline iterations (Section 4.2). We call this dataset $D_{R}$. We intentionally select these “easier” CVEs because they serve as a controlled probe for ablation: if performance degrades even on CVEs that CVE-Genie can normally handle reliably, it indicates that the missing information is truly critical. As shown earlier, the availability of the PoC is significant. To simulate other incomplete context scenarios we create three variants of $D_{R}$ by systematically removing: (a) the security advisory, (b) the patch commit, and (c) all except the CVE description.
@@
Methodology. For each variant of $D_{R}$, we run CVE-Genie three times per CVE.
Results. Removing the patch commit lowers performance to 10/15 (67%), even when the PoC is available, as it significantly impairs the model’s ability to localize the vulnerability. For example, in CVE-2025-31481, an authorization flaw in the PHP-based API Platform Core, the system lacked direction, explored irrelevant parts of the codebase, and eventually exhausted its cost budget. Removing the security advisory also results in a drop to 11/15 (73%). Most failures here are due to environment setup and verification issues rather than exploit synthesis errors. As security advisories often contain contextual details that guide environment construction and verification strategy. On the other hand, even with only the CVE description, CVE-Genie successfully reproduces 9/15 (60%), particularly for vulnerabilities with recognizable patterns (e.g., SSRF, CSRF). This highlights the model’s baseline robustness under minimal context.
RQ2: CVE-Genie can reproduce CVEs with limited information. Code-based PoCs are most helpful, but in their absence, patch commits offer valuable localization cues. However, reproduction success decreases with minimal context.
4.4 Large Scale Evaluation
@@
Dataset $(D_{L})$. For our large-scale evaluation, we curate a dataset, $(D_{L})$, comprising all CVEs published between June 2024 and May 2025 that include publicly available source code. Since these CVEs postdate knowledge cutoff of LLMs (o3 and o4-mini) used in CVE-Genie, they ensure zero prior exposure. $D_{L}$ includes 841 CVEs across 186 CWEs, 29 programming languages, and 440 open-source projects. Our statistical analysis shows approximately 55% of CVEs in $D_{L}$ fall under the top 25 most dangerous CWEs [37], and around 49% lack a PoC in their advisory.
Results. CVE-Genie reproduced 428 CVEs out of 841, spanning 22 programming languages, 141 vulnerability types, and 267 projects (see Table 4). To better understand outcomes, we manually analyzed 50 successful and 50 failed cases, and categorize the main reasons for successes and failures.
Successes. Web vulnerabilities (XSS, CSRF, SQLi, path traversal) were most reproducible, especially in interpreted languages (Python, JavaScript, PHP, Ruby, TypeScript), as shown in Table 4. Projects with clear setup instructions and code-based PoCs with crashing input had the highest success rates, as runnable PoCs directly guided exploit construction.
Failures. Memory-safety bugs (C/C++), concurrency issues, and UI-dependent flaws were hardest to verify. Many failures were due to brittle builds, poor project documentation, or resource-heavy environments; however, 24% failed from time/cost overruns. PoCs expressed only in natural language contributed little, often leaving too much ambiguity for reliable reproduction. Compiled toolchains (C/C++, Rust, Go) were prone to the most setup problems.

### v2

@@ 3.5.1 Builder
In this section, we systematically select the most optimal configuration of LLMs and prompts for the Builder.
@@ 3.5.2 Exploiter
In this section, we systematically select the most optimal configuration of LLMs and prompts for the Exploiter.
@@
Candidate Selection. Generating a working exploit from a given CVE in a pre-configured vulnerable environment is a crucial task for reproducing CVEs. However, the Exploiter requires less code base exploration than the Builder, as advisories often provide PoCs, exploitation steps, or vulnerability details. Thus, we include expensive reasoning models like o3, claude-3.7-sonnet, and gemini-2.5-pro in the candidates list of developers. To mitigate instability seen in open-source models (Section 3.5.1), we employ only closed-source LLMs as developers, while using the same critic LLMs as the Builder.
Developer Evaluation. For the 32 successful setups from Section 3.5.1, we ran CVE-Genie up to the Exploit Developer stage 6 and manually evaluated each developer LLM as follows: (1) we checked whether the LLM executed a PoC or command sequence that demonstrated the exploit, (2) we verified that the vulnerability was clearly triggered (e.g., uncontrolled recursion in CWE-674 resulting in a RecursionError), (3) we assessed fidelity to the CVE description, (4) we ensured no mock-up environments or fake exploits were used, and (5) we confirmed the exploit script followed the expected format (see Section 3.3.1). Among all models, o3 consistently performed best, successfully reproducing the highest number of exploits (13 out of 27) and demonstrating them end-to-end within the actual vulnerable environment. In contrast, o4-mini often produced incorrect or mock-up exploits and struggled with runtime errors. While claude and gemini models showed strong capabilities in demonstrating exploits, they lacked versatility across different vulnerability types. Hence, we select o3 as the Exploit Developer.
Critic Prompt Design. We examined the failed exploit attempts and categorized the recurring error patterns, which directly mirrored our manual verification criteria (e.g., lack of exploit demonstration, unclear vulnerability trigger, deviation from CVE description, reliance on dummy/fake exploits, or incorrect script formatting). We then incorporated these categories into the critic agent’s prompt, enabling it to automatically detect these issues for Exploit Developer.
@@
Feedback Assessment. We evaluated feedback effectiveness by running CVE-Genie on 32 correct setups with the top developer (o3) and critic (o4-mini). This yielded valid exploits for 16 CVEs, of which 3 were fixed based on the critic’s feedback. As in Section 3.5.1, the developer mainly resolved minor issues (e.g., incomplete verification or missing log triggers), while more complex flaws (e.g., incorrect exploit verification) persisted. To balance usefulness and cost, we therefore restrict feedback to a single iteration.
@@ 3.5.3 Verifier
In this section, we systematically select the most optimal configuration of LLMs and prompts for the CTF Verifier.
@@
Critic Evaluation. As shown in Figure 7(b), only the o3 model successfully identified critical but subtle issues in the verifier scripts, such as weak or flawed verification logic (Figure 4). Other models provided weak or no critique, making o3 the most
Feedback Assessment. We evaluated CVE-Genie on 16 correct exploits from the Exploiter, successfully generating accurate verifiers for 15 of them, of which 4 verifiers were improved through the critic’s feedback. In our experiment, the maximum number of feedback iterations required to resolve either functional or critique-based issues was 4; therefore, we allow up to 5 retries for both the Flag Checker and the Verifier Critic.
@@ 3.5.4 Time and Cost Constraints
4 Experimental Methodology
In this section, we conduct a four-part evaluation of CVE-Genie: (1) a baseline study assessing consistency and efficiency (Section 4.1); (2) an ablation study to assess the impact of key components of CVE-Genie on its scalability and effectiveness (Section 4.2); (3) robustness testing under incomplete CVE information (Section 4.3); and (4) large-scale evaluation on 841 CVEs (Section 4.4). These evaluations are guided by the following research questions:
  • RQ1: Does the performance of CVE-Genie vary over multiple runs of CVE reproduction?
  • RQ2: Is the complex architecture of CVE-Genie necessary, or can standalone LLMs achieve comparable results?
  • RQ3: Can CVE-Genie effectively reproduce CVEs with limited information, and which CVE report components are most crucial for reproduction?
  • RQ4: Can CVE-Genie reproduce a broad range of CVEs across diverse CWEs, languages, and projects?
4.1 Baseline Assessment
In this section, we evaluate the baseline performance of CVE-Genie and examine how its reproduction success evolves across multiple runs in the presence of LLM non-determinism.
Dataset. For this study we use the same $D_{S}$ dataset (Section 3.5).
Methodology. Due to the non-deterministic nature of LLMs, a single run may fail due to incorrect reasoning, while later attempts might succeed [33]. To account for this, we run CVE-Genie multiple times on $D_{S}$, using its most optimal configuration (Section 3.5). After each iteration, we store successfully reproduced CVEs and rerun on the remaining ones. This iterative approach helps maximize CVE reproduction and reveals CVE-Genie’s convergence.
Results. We perform manual analysis of randomly selected 25 CVE reproduction runs, and observe that the variability in CVE-Genie’s performance primarily stems from the Builder phase, which is more open-ended than the targeted exploit generation guided by CVE context. Project setup involves repository exploration (Pre-Requisite Developer) and command execution (Setup Developer), where environment-specific behaviors often cause failures. For example, one run failed because the Setup Developer modified the PATH variable, while a clean retry succeeded, highlighting the value of reattempting in a fresh environment. Since the CTF Verifier depends on the Exploiter, which in turn relies on the Builder, this initial project setup is a critical bottleneck. As a previously failed build, if later successful, can enable downstream modules to reproduce a new CVE (as illustrated in Figure 8). Over multiple runs, the number of successful project, exploit, and verifier builds showed consistent gains with convergence at run seven, indicating that most recoverable failures can be resolved within a few retries.
RQ1: Due to the non-determinism of LLMs, CVE-Genie exhibits variability across runs, primarily during project setup and environment configuration. Once project setup succeeds, downstream exploit generation and verification are more stable and converge in a few retries.
@@ 4.1 Baseline Assessment
4.2 Importance of CVE-Genie’s Architecture
Having evaluated the selection of LLMs and prompts for CVE-Genie (Section 3.5), we now assess CVE-Genie’s architecture via a systematic ablation study that isolates the impact of CVE-Genie’s individual components on end-to-end CVE reproduction.
Dataset. We perform this ablation study using 15 CVEs from $D_{S}$, each with complete information (i.e., CVE description, PoC, patch commit, and security advisory) and reproduced consistently in the first three iterations of baseline assessment (Section 4.1), and call this dataset $D_{R}$. By keeping the CVE context complete, we ensure that any performance differences arise solely from modifications
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    Single Monolithic Agent: Combine all modular agents into a single agent with access to the same tools, thereby evaluating the performance of a standalone LLM without CVE-Genie’s agentic design and structured guidance (see Appendix G).
Results. In our ablation study (Table 4), removing the Knowledge Builder component reduces reproduction success to 9/15, as agents struggle to interpret unstructured advisory data and hit context window limits more often, which results in a 30% increase in exploit failures. Eliminating the Pre-Requisite Developer has a comparatively mild effect (13/15), but places additional burden on the Setup Developer as it has to do both codebase exploration and project setup at the same time, leading to 27% more tool-call limit violations. We observe this behavior particularly in complex repositories such as CVE-2025-32389 in Nameless, a website software for Minecraft servers, with not very clear setup guides. Feedback loops emerged as the most critical design element, and without them the reproduction success dropped sharply to 5/15 (–67%). Similarly, removing critic agents reduced success to 8/15 while increasing false reproductions by 47%, highlighting their role in reliable end-to-end CVE reproduction. Finally, collapsing all components into a single monolithic agent resulted in zero successful reproductions, even when using advanced standalone LLMs (o3), underscoring the necessity of CVE-Genie’s modular, feedback-driven, and critic-guided
RQ2: Standalone LLMs are currently unable to reproduce CVEs end-to-end. CVE-Genie’s modular, feedback-driven, and critic-guided architecture design is essential for reliable CVE reproduction, and removing any of its components significantly degrades performance (Table 4).
Ideal Case Knw Build Pre- Reqs Feed-back Critic Agents Single Agent
15 / 15 9 / 15 13 / 15 5 / 15 8 / 15 0 / 15
Table 4: Ablation study of CVE-Genie’s design.
@@ 4.3 Robustness to Loss of CVE Context
To evaluate how CVE-Genie performs when critical components of a CVE report are missing, we study its robustness under systematically reduced CVE context. We first present a detailed qualitative case study of a single CVE to illustrate how CVE-Genie adapts its reasoning under different context losses, and then conduct a quantitative ablation study to quantify the impact at scale.
4.3.1 Case Study for CVE-2024-4340.
We begin by performing a detailed case study for CVE-2024-4340 reproduction under three levels of CVE context losses, and compare their reproductions against the ideal setting shown in Figure 2.
1) No Security Advisory / PoC. In this setting, all security advisories are removed from the CVE context, leaving only the developer patch commit [^2] CVE-2024-4340 Developer Patch: https://github.com/andialbrecht/sqlparse/commit/b4a39d9850969b4e1d6940d32094ee0b42a2cf03 and the CVE description [^3] CVE-2024-4340 CVE Description: “Passing a heavily nested list to sqlparse.parse() leads to a Denial of Service due to RecursionError.”. As a result, CVE-Genie relies heavily on analyzing the patch commit, exploring modified files, and inferring the PoC from the regression test added by the developer in tests/test_regressions.py. This test closely resembles the PoC described in the existing advisory [^4] CVE-2024-4340 GHSA: https://github.com/advisories/GHSA-2m57-hf25-phgg, and the generated exploit payload (Figure 9(a)) is comparable to the ideal run shown in Figure 3(a). Due to the absence of explicit PoC information, this run required additional exploration, resulting in a longer execution time (9 minutes) and higher cost (\$0.67) compared to the ideal scenario.
2) No Patch. Here, the developer patch commit is removed while security advisories remain available. CVE-Genie primarily leverages the PoC provided in the advisory, which includes clear PoC exploit code, execution traces, vulnerable code, and vulnerability root cause. Consequently, the reproduction closely matches the ideal run in terms of generated PoC exploit (Figure 9(a)), verification scripts, execution time, and cost.
3) Only CVE Description. In the most constrained setting, CVE-Genie has access only to the CVE description. Based on this limited information, it starts by identifying the implementation of sqlparse.parse and hypothesizes on how a deeply nested inputs can trigger a recursive execution path leading to a RecursionError. During PoC exploit generation and verification, the first run failed to produce a payload that successfully triggered the RecursionError. However, a second run from scratch succeeded in generating a valid payload (shown in Figure 9(b)). This setting proved to be the most challenging, requiring 20 minutes of execution time and \$4.50 LLM cost over two runs.
1payload = ’[’ * depth + ’]’ * depth
2sqlparse.parse(payload)
(a) Security advisory/PoC or patch commit is not available.
1payload = ’select’ + ’(’ * depth + ’1’ + ’)’ * depth
2sqlparse.parse(payload)
(b) Only CVE description is available.
Figure 9: PoC exploit payload generated by CVE-Genie for CVE-2024-4340 under varying levels of incomplete CVE context compared to the payload in ideal scenario in Figure 3(a).
4.3.2 Evaluation at Scale.
In this section, we perform an ablation study to systematically quantify the robustness of CVE-Genie to incomplete CVE context at scale.
@@
Dataset ($D_{R}$). We perform this ablation with $D_{R}$ dataset (Section 4.2). We intentionally select these “easier” CVEs as a controlled probe, since CVE-Genie can consistently reproduce them under full context. Hence, any performance degradation after removing specific information indicates that the missing context is genuinely critical. To simulate incomplete CVE context, we construct three variants of $D_{R}$ by systematically removing: (a) security advisory/PoC, (b) patch commit, and (c) everything except CVE description.
@@
Methodology. In this study, for each variant of $D_{R}$, we run CVE-Genie three times per CVE, iteratively.
Results. Removing the developer patch commit leads to a notable drop in reproduction success 10/15 (67%), even when a PoC is available. Patch commits provide essential localization signals by identifying relevant files and code paths. Without patch commits, the model must rely on broad codebase exploration, often resulting in inefficient search and budget exhaustion. We observed this behavior in CVE-2025-31481 reproduction, an authorization flaw in the PHP-based API Platform Core, where the absence of a patch caused CVE-Genie to explore unrelated components and fail to converge. Excluding security advisories reduces the success rate to 11/15 (73%). In these cases, failures are primarily attributable to environment setup and verification difficulties rather than exploit generation. Advisories often provide crucial contextual information, such as dependency versions, execution constraints, and validation guidance, that streamline reproduction, particularly when PoCs are not provided as executable code but rather as high-level textual descriptions. Despite these challenges, CVE-Genie maintains baseline robustness under minimal context. When only the CVE description is provided, it successfully reproduces 9/15 cases (60%), albeit with increased cost and runtime. This demonstrates that while comprehensive CVE reports significantly improve efficiency and reliability, meaningful reproduction remains feasible even
RQ3: CVE-Genie can reproduce CVEs under limited context. Code-based PoCs provide the strongest guidance, but in their absence, patch commits help localize vulnerabilities. Reproduction remains possible even with only CVE descriptions, but with lower success rates, higher cost, and increased time.
@@
Project Type CVEs Reprod. Projects Reprod.
AI/ML Platform 22/38 (57.9%) 14/16 (87.5%)
Web Application/Backend 171/383 (44.6%) 86/156 (55.1%)
Library/Framework 163/224 (72.8%) 115/147 (78.2%)
Cloud/DevOps Application 13/33 (39.4%) 13/25 (52.0%)
Operating System/Runtime 15/31 (48.4%) 5/15 (33.3%)
Desktop Application 6/34 (17.6%) 2/13 (15.4%)
CLI Tool/Utility 18/37 (48.6%) 15/28 (53.6%)
Blockchain/Crypto/FinTech 1/13 (7.7%) 1/7 (14.3%)
Mobile Application/SDK 2/7 (28.6%) 2/7 (28.6%)
Embedded/Networking 11/27 (40.7%) 10/18 (55.6%)
Security Tool/Server 6/12 (50.0%) 4/6 (66.7%)
Table 5: CVEs reproduction coverage by project type.
CWE-ID Overall success PoC available No PoC
CWE-79 65/142 (45.8%) 43/94 (45.7%) 22/48 (45.8%)
CWE-200 20/43 (46.5%) 15/20 (75.0%) 5/23 (21.7%)
CWE-22 16/37 (43.2%) 11/22 (50.0%) 5/15 (33.3%)
CWE-284 19/36 (52.8%) 16/26 (61.5%) 3/10 (30.0%)
CWE-400 22/36 (61.1%) 13/19 (68.4%) 9/17 (52.9%)
CWE-20 13/29 (44.8%) 6/15 (40.0%) 7/14 (50.0%)
CWE-770 14/24 (58.3%) 7/9 (77.8%) 7/15 (46.7%)
CWE-918 7/22 (31.8%) 6/13 (46.2%) 1/9 (11.1%)
CWE-89 11/21 (52.4%) 8/15 (53.3%) 3/6 (50.0%)
CWE-1333 15/19 (78.9%) 8/10 (80.0%) 7/9 (77.8%)
CWE-94 9/19 (47.4%) 2/8 (25.0%) 7/11 (63.6%)
CWE-863 7/15 (46.7%) 4/6 (66.7%) 3/9 (33.3%)
CWE-287 8/15 (53.3%) 4/4 (100.0%) 4/11 (36.4%)
CWE-285 7/15 (46.7%) 5/8 (62.5%) 2/7 (28.6%)
CWE-269 7/13 (53.8%) 4/8 (50.0%) 3/5 (60.0%)
CWE-122 8/13 (61.5%) 2/4 (50.0%) 6/9 (66.7%)
CWE-502 6/13 (46.2%) 3/6 (50.0%) 3/7 (42.9%)
CWE-532 6/12 (50.0%) 2/2 (100.0%) 4/10 (40.0%)
CWE-78 4/11 (36.4%) 4/9 (44.4%) 0/2 (0.0%)
CWE-352 8/10 (80.0%) 5/6 (83.3%) 3/4 (75.0%)
CWE-74 7/10 (70.0%) 4/4 (100.0%) 3/6 (50.0%)
CWE-416 3/10 (30.0%) 0/0 (–) 3/10 (30.0%)
CWE-639 2/9 (22.2%) 1/4 (25.0%) 1/5 (20.0%)
CWE-116 4/8 (50.0%) 1/3 (33.3%) 3/5 (60.0%)
CWE-347 7/8 (87.5%) 4/4 (100.0%) 3/4 (75.0%)
Table 6: Top 25 CWEs identified in the large-scale study and their reproduction success rates, overall and stratified by PoC availability.
@@
4.4 Large Scale Evaluation
Dataset $(D_{L})$. For our large-scale evaluation, we construct a dataset, $(D_{L})$, consisting of all CVEs published between June 2024 and May 2025 that include publicly available source code. A CVE is included in $(D_{L})$ if at least one of the following conditions is satisfied: (1) the CVE entry contains a GitHub URL (e.g., a patch commit) in its references that links directly to the relevant source code, or (2) the CVE specifies affected software versions, and the source code corresponding to those versions is publicly accessible. Because these CVEs were disclosed after the knowledge cutoff of the LLMs (o3 and o4-mini) used in CVE-Genie, $(D_{L})$ ensures zero prior exposure during model training. Overall, $(D_{L})$ comprises 841 CVEs spanning 186 CWEs, 29 programming languages, and 440 distinct open-source software (OSS) projects. The OSS projects cover a broad range of application domains, including web applications and backends (156), libraries and frameworks (147), CLI tools and binaries (28), cloud and DevOps systems (25), embedded and networking software (18), operating systems and runtimes (17), AI/ML platforms (16), desktop applications (13), blockchain and cryptocurrency systems (7), mobile applications and SDKs (7), and security tools (6). In terms of vulnerability types, the majority of CVEs in $(D_{L})$ fall under the following categories: injection and improper neutralization (29%), authorization and authentication (11%), resource management (8%), information exposure (6%), memory safety (3%), request forgery (3%), and others. Our statistical analysis further shows that approximately 55% of CVEs in $D_{L}$ fall under the top 25 most dangerous CWEs [37], and around 49% lack a PoC in their advisory.
Results. CVE-Genie reproduced 428 CVEs out of 841 (Figure 10), spanning 22 programming languages, 141 vulnerability types, and 267 projects. To better understand outcomes, we perform statistical analysis on all reproduction runs as well as manually analyze 50 successful and 50 failed cases, and categorize results as follows:
Domain-Dependent Effectiveness of CVE-Genie. We observe that CVE-Genie shows the highest performance for library frameworks and AI/ML platforms categories, with the reproduction rates of 78% and 58%, excelling on CWEs such as CWE-79 (XSS), CWE-89 (SQLi), and CWE-284 (access control). Meanwhile, the success appears more moderate for web/backend applications (45%), where logic-driven CWEs such as 200, 22, and 918 have lower success rates. CVE-Genie struggles in binary-heavy domains such as desktop (18%), blockchain (8%), and mobile (29%) applications, especially with memory-safety (e.g., CWE-122, 416) and crypto/validation (CWE-295, 327) flaws. A finer-grained breakdown by vulnerability type and project characteristics is provided in the Appendix, where Table 6 reports reproduction success rates for the top CWEs. Meanwhile, Table 7 and Table 5 further contextualize results by project scale, language, and domain. Finally, Tables 919 detail domain-specific CWE distributions, illustrating how failure
Failures. For $198$ cost/time overrun failures, ~80% were due to cost overruns, where the agent either got stuck during project builds (~41%) or while generating exploits (~59%). The majority of failures involved web application CVEs lacking either vulnerability details or setup documentation. For example, none of the CVEs in project discourse built successfully since its README only links to external documentation, which our agents cannot access, causing the exhaustion of their budgets in exploring the codebase.